OpenStudy (anonymous):

Solve for the variable: 3w= 4 1/5 A. -7/5 B. 7/5 C. 63/5 D. -63/5

OpenStudy (anonymous):

Well whats the equation?

OpenStudy (anonymous):

divide both sides by 3

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

63/5

OpenStudy (texaschic101):

3w = 4 1/5 3w = 21/5 -- divide by 3 w = (21/5)/3 w = 21/5 * 1/3 w = 21/15 or 1 6/15

OpenStudy (anonymous):

I got 7/5 @texaschic101

OpenStudy (texaschic101):

7/5 is the same as 21/15

OpenStudy (texaschic101):

your right.....I didn't reduce